OpenStudy (anonymous):

Write an equation in slope-intercept form of the line satisfying the given conditions. Through (-7, 4), parallel to 7x + 5y = -19 Would it be y=7/5x+29/5 or y=-5/7x+4/7 ? > _ <

OpenStudy (anonymous):

y = -7/5 - 3.8 I think

OpenStudy (anonymous):

7x+5y=-19 -7,4 5y=-7x-19 y=- 7x / 5- 19 / 5 m=- 7 / 5 y=mx+b 4=(-7 / 5 )*(-7)+b b=-(29)/(5) y= - 7x / 5 - 29 / 5

OpenStudy (anonymous):

7x + 5y = -19 5y = -7x -19 y = -7/5x - 19/5 (-7,4) y2-y1=m (x2-x1) y-4 = -7/5 (x--7) y-4=-7/5(x+7) y-4= -7/5x - 49/5 y = -7/5x - 49/5 + 4 y = -7/5x -29/5 this is the answer I believe

OpenStudy (anonymous):

I got the same answer as @Luis_Rivera so it is indeed correct.

OpenStudy (anonymous):

Thank you so much c: I'm in the middle of an assignment so I have a few more coming > _ >'